OpenStudy (anonymous):

Jason weighs 36 pounds. Together, he and his sister weigh four pounds more than three times the weight of his sister. What is the weight, w, of Jason’s sister? 36 + w = 3w; 18 pounds 36 + w – 4 = 3w; 19 pounds 36 + w = 3w + 4; 16 pounds 36 + w = 3w – 4; 20 pounds

OpenStudy (anonymous):

ergh Jason weighs 36 pounds. Together, he and his sister weigh four pounds more than three times the weight of his sister. What is the weight, w, of Jason's sister? (1 point) 36 + w = 3w; 18 pounds 36 + w – 4 = 3w; 19 pounds 36 + w = 3w + 4; 16 pounds 36 + w = 3w – 4; 20 pounds

OpenStudy (anonymous):

36+w=3w +4; 16 pounds

OpenStudy (anonymous):

36 + w = 3w; 18 pounds 36 + w - 4 = 3w; 19 pounds 36 + w = 3w + 4; 16 pounds 36 + w = 3w - 4; 20 pounds

OpenStudy (anonymous):

really? thanks. i have a few more @heavenfun13 can you help with those too?

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

Labrina received a $50 gift card to an online store. She wants to purchase some bracelets that cost $8 each. There will be a $10 overnight shipping fee. How many bracelets can she buy? 4 bracelets 5 bracelets 6 bracelets 7 bracelets

OpenStudy (anonymous):

i think its 5

OpenStudy (anonymous):

ok let me work it out

OpenStudy (anonymous):

5 its right

OpenStudy (anonymous):

Bobby hopes that he will someday be more than 70 inches tall. He is currently 61 inches tall. How many more inches, x, does Bobby need to grow to reach his desired height?

OpenStudy (anonymous):

i need to equation for this one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

61+x or 16x